Solve the following differential equation with given initial conditions using the Laplace transform. y + 5y + 6y = ut - 1)
Answer #1

Q. 1. Answers Given y try toy u(t-1)-8 (+-2) f yo) = -2 y los-5 Taking Laplace on both the sides. Layile +5L{ys to Lf yg =For sim Sinflicting lets write you sy oo s?y +s 5 +5( SY +2 +6y 25 s +3+58 +6y +38-5+10 es - e25 -> (3P+58 +6 Y . (@s+5) s -now Ls (3+2) (1+2) = tē s 41 5 6 % s 2 st2 3 8+3 6 s 27 st. St3 where Note:- Iseas.pl = ft-a).ulta) fres = 1 1 1 . Ut-1)] - 1Thow 16+2363+3) str 843 st2 ees By ett-2). 44-2 St2 843 e3lt-2). ust-2) e finally, let 24+5 (4+2)(8+3) - A &+2 4B St3 => 28+5AO Stron Gist ਨ+] +2) +3) | stol olt test () o y = es + 4 (813 ) ਨੂੰ+91)9 (343 ) how, taking inverse Laplace L - Y & = ੫੬ (18now uhrg evu get पण =148) मो.-) K 2 हस्.िuff- हरविले. परिण t-u-D -हम- Answer.